Paul is employed as an auditor by a CPA firm. On most days, he commutes by...

  1. Paul is employed as an auditor by a CPA firm. On most days, he commutes by auto from his home to the office.

During one month, however, he has an extensive audit assignment closer to home. For this engagement, Paul drives directly from home to the client’s premises and back. Mileage information is summarized below:

Home to office 12 miles

Office to audit client 15 miles

Audit client to home 10 miles

If Paul spends 20 days on the audit and provides an adequate accounting to his employer, what is his reimburseable mileage?
Answer #1

As Paul directly drive from home to the client premises the total miles travelled by him will be calculated taking into consideration the distance between the audit client to home i.e 10 miles per side

Total miles travelled by Paul in 20 days i.e. 40 sides = 40*10 = 400 miles

Therefore is reimbursable mileage will be 400 miles

As per internal revenue service standard mileage rate for year 2020 is 57.5 cents per mile

So total reimbursement will be = 57.5*400 = 23,000 cents = $ 230